OpenStudy (anonymous):

next question: 16^x A. 4*(4^x) B. 4*(4^2x) C. 4^2*4^x D. 4^(2x) E. 4^x*4^x F. (4*4)^x

OpenStudy (anonymous):

which expressions are equivalent to the one given above? select all that apply.

hero (hero):

Hint: 4 x 4 = 16

OpenStudy (anonymous):

I said E&F so far

OpenStudy (anonymous):

are those the only answers

OpenStudy (anonymous):

ok

hero (hero):

Wait, what abotut D?

OpenStudy (anonymous):

& then what abt this problem?

OpenStudy (anonymous):

wait what abt d?

hero (hero):

Think about it

OpenStudy (anonymous):

blurbendy was saying the same ting & the he asked me to get a 2nd opinion on it that's why I asked u

OpenStudy (anonymous):

wait but how could d be a possiblity

hero (hero):

Double check your rules of exponents

OpenStudy (anonymous):

WAIT! r u saying that D is also an answer???

hero (hero):

\[4^{2x} = (4^2)^x = 16^x\]

OpenStudy (anonymous):

shoot! r u serious! that was the one I corrected and I already submitted it!

OpenStudy (anonymous):

ugh! why didn't I think of that. that was a silly mistake!!!

hero (hero):

Why did you wait until after you submitted it to check?

OpenStudy (anonymous):

no, its kind of like a timed thing.... if you wait too long the screen changes and goes on the next question

hero (hero):

Well you need to be faster. How long do you get per question?

hero (hero):

And where's your calc? Use that

OpenStudy (anonymous):

its not like a specific timing

hero (hero):

How do you know the time isn't specific?

hero (hero):

You have to time it yourself probably.

OpenStudy (anonymous):

i just know that once you take too long (maybe more than like 5 mins or something) it just t let u er the question

hero (hero):

Time it with a question you already know the answer to.

hero (hero):

5 minutes is plenty of time.

OpenStudy (anonymous):

ok well here i want to check the next one w/ u

OpenStudy (anonymous):

same directions

hero (hero):

Sounds like a test though.

OpenStudy (anonymous):

3^x A. 3*(3^(x+1)) B. 3*(3^(x-1)) C. (18^x)/3 D. (18/6) E. x^3 F. (18^x/6^x)

OpenStudy (anonymous):

so i got..... D, F,

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

I just need help w/ the question I asked right now

OpenStudy (anonymous):

the above comments are abt another question

OpenStudy (anonymous):

for this question I got B, D, & F. Is that correct?

mathslover (mathslover):

I can write : \(3^x\) as \(3^x\times \frac{3}{3}\) Or , \(\cfrac{3^x}{3} \times 3\) \(3^{x-1} \times 3\)

OpenStudy (anonymous):

so is my answer correct?

mathslover (mathslover):

Yeah it is B) only.

OpenStudy (anonymous):

how is it b ONLY? that doesn't really make sense. 18/6 equals 3, so 3^x wold be correct, right?

OpenStudy (anonymous):

same thing with D,?

mathslover (mathslover):

18/6 = 3 \(\ne\) 3^x

mathslover (mathslover):

3 is not equal to 3^x.

OpenStudy (anonymous):

yes but it raised to x

mathslover (mathslover):

3 is 3^1 while, 3^x

OpenStudy (anonymous):

so wouldn't it be correct

mathslover (mathslover):

x is an variable, it can eitherr be 0, 1,2,3 or anything. So, only B) is correct

mathslover (mathslover):

F is also Right,

mathslover (mathslover):

\(\cfrac{18^x}{6^x } = 3^x\)

OpenStudy (anonymous):

yea, so then it'd be B, D & F correct?

mathslover (mathslover):

No. B and F

OpenStudy (anonymous):

but D is the same thing as F....

mathslover (mathslover):

No, both are different. 18^x / 6^x gives 3^x and 18/6 gives 3 . 3^x and 3 are completely Different.

OpenStudy (anonymous):

because the product law states that: example: (2*8)^3 is the same thing as (2^3)*(8^3)

OpenStudy (anonymous):

are you like still thinking

OpenStudy (anonymous):

r u there :( ?

OpenStudy (anonymous):

:( :(

hero (hero):

No

OpenStudy (anonymous):

wut?

OpenStudy (anonymous):

next question: 16^x A. 4*(4^x) B. 4*(4^2x) C. 4^2*4^x D. 4^(2x) E. 4^x*4^x F. (4*4)^x

OpenStudy (anonymous):

I said B, D, & F. I'm wondering if that's correct

hero (hero):

Bro, we already went over this question before. I think there's something wrong with you.

hero (hero):

@music101

hero (hero):

And you don't even remember the correct answer choices... @music101

OpenStudy (anonymous):

just mind ur own business..... this question's already closed...

hero (hero):

Originally, you picked E and F. Then I pointed out that D was also correct.

hero (hero):

Wow bro....

OpenStudy (anonymous):

stop calling me that ok.... leave me alone..... I was just double-checking. Sheesh!

OpenStudy (anonymous):

sorry if u think im being rude. its called a closed question though

hero (hero):

Review your rules